Tripcurl Surfboards manufactures fiberglass surfboards. The standard cost of direct materials and direct manufacturing labour is.100 dollars per board. This includes 20 pounds of direct materials, at the budget prices of $2 per pound and 5hours of direct manufacturing labour, at the budget rate of $12 per hour. Following our additional data for the month of July: Units completed Direct material purchases Cost of direct material purchases Actual direct manufacturing labor hours32,000 hours Actual direct labor cost 6,000 units 150,000 pounds $292,500 $368,000 126,250 pounds Actual direct material used There are no beginning inventories. Required: 1. Calculate direct material price (based on purchases) and efficiency variances for July. 2. Calculate direct manufacturing labor price and efficiency variances for July. 3. Tripcurl Surfboards is considering using variances calculated in parts 1) and 2) to evaluate the performance of its materials purchasing manager and production manager Discuss any two limitations which the business should be aware of when using variances to evaluate manager performance.
Answer #1

given data

units completed = 6000 units

direct material purchased = 150000 pounds

cost of direct material purchased = 292500

actual direct material used = 126250 pounds

actual direct manufacturing labor hours = 32000 hours

actual direct labor cost = 368000

standaed price of materials = $ 2 per pound

1) calculate direct material price varience and efficiency variency :

actual price of material = cost of direct material purchased / direct material purchased

= 292500 / 150000

= 1.95

direct material price varience = direct material purchased (standard price - actual price)

= 150000 * (2 - 1.95)

= $ 7500

standard quantity for actual production = units completed * cost of direct materials

= 6000 * 20

= 120000

direct material efficiency variances = standard price * (standard quantity for actualproduction -actual quantity)

= 2 * (120000 - 126250)

= (12500)

2) calculate direct manufacturing labor price and efficiency varience :

actual price of labor = actual direct labor cost / actual direct manufacturing labor hours

= 368000 / 32000

= $ 11.5

labor price varience = actual hours worked * (standard price - actual price)

= 32000 * (12 - 11.5)

= $ 16000

standard hours for actual production = units completed * direct manufacturing lobor rate

= 6000 * 5

= 30000 hours

labor efficiency varience = standard price per hour * (standard hours for actual production - actual hours)

= 12 * (30000 - 32000)

= (24000)

3) limitations :

performance of its materials purchasing manager :

  • using the quality (lower /higher) of materials when the importance of the production
  • unexpected changes in market rates of materials by inflation or defflation
  • quantity discounts will be given for attraction of customers

performance of its materials production manager :

  • unexpected changes in production techniques
  • maintain an effective lobor for better performance
  • increase labor when work loaded, decrese labor when free time of production
  • some unexpected defect in machinary which could result in wastage of material.